OpenStudy (anonymous):

1. Is (2,4) a solution to the equation y = x – 2? 2. Is (-3,-9) a solution to the equation y = 3x? 3. Which of (3, 5), (4, 6), (5, 7), and (6, 8) are solutions to y = x + 2?

OpenStudy (anonymous):

The best way to handle this is to simply plug the (x,y) values into the equation and verify that the equality is met. (i.e): y=x-2 plugging in (2,4) means that 4=(2-2), which is obviously a false statement (i.e. 4=0 is false)

OpenStudy (anonymous):

well all are similar questions. For example, the first one: Is (2,4) a solution to y=x-2 ? What we do, is plug in the values x=2 and y=4 into the equation and see if it's true or not y=x-2 --> replacing x and y with their values, we get: (4)=(2)-2 Which means 4=0 And this is not true! So, (2,4) is NOT a solution for y=x-2

OpenStudy (anonymous):

thanks Guys

OpenStudy (anonymous):

what about the other two questions ?

OpenStudy (anonymous):

can you try them and we'll check your work ? or let us know where you have difficulty

OpenStudy (anonymous):

i did em okay for number two i said its not a solution & number three i said ( 4,6 ) & ( 5.7 ) am i correct ?